# Learned Hemodynamic Coupling Inference in Resting-State Functional MRI

### v1

@@ Abstract
Functional magnetic resonance imaging (fMRI) provides an indirect measurement of neuronal activity via hemodynamic responses that vary across brain regions and individuals. Ignoring this hemodynamic variability can bias downstream connectivity estimates. Furthermore, the hemodynamic parameters themselves may serve as important imaging biomarkers. Estimating spatially varying hemodynamics from resting-state fMRI (rsfMRI) is therefore an important but challenging blind inverse problem, since both the latent neural activity and the hemodynamic coupling are unknown. In this work, we propose a methodology for inferring hemodynamic coupling on the cortical surface from rsfMRI. Our approach avoids the highly unstable joint recovery of neural activity and hemodynamics by marginalizing out the latent neural signal and basing inference on the resulting marginal likelihood. To enable scalable, high-resolution estimation, we employ a deep neural network combined with conditional normalizing flows to accurately approximate this intractable marginal likelihood, while enforcing spatial coherence through priors defined on the cortical surface that admit sparse representations. The proposed approach is extensively validated using synthetic data and real fMRI datasets, demonstrating clear improvements over current methods for hemodynamic estimation and downstream connectivity analysis.
I Introduction
I-A Motivation
@@
Functional magnetic resonance imaging (fMRI) is a non-invasive imaging technique widely used to study neural activity. fMRI collects time-dependent blood oxygenation level–dependent (BOLD) signals, which reflect hemodynamic changes associated with neural activity [1]. Accurate characterization of the hemodynamic coupling, i.e., the forward model that maps neural activity to BOLD signals, enables inference of underlying neural activity by solving the hemodynamic inverse problem [2]. However, this problem is complicated by the fact that the hemodynamic coupling is partially unknown and varies both across brain regions and between individuals [3, 4]. Failure to account for the spatially varying hemodynamics can bias downstream estimates of functional connectivity [5, 6, 7]. Additionally, the hemodynamic properties are themselves of independent interest as potential imaging biomarkers. Indeed, systematic hemodynamic alterations have been identified in a variety of neuropsychiatric conditions, including autism [8], schizophrenia and bipolar disorder [9], and obsessive compulsive disorder [10].
@@
There exists a substantial body of work on estimating hemodynamic coupling in task/event-based fMRI designs [11, 12, 13, 14]. In this setting, the neural signals are assumed to be known from the experimental design, making inversion considerably easier than in resting-state fMRI (rsfMRI), where both neural signals and hemodynamic response are unknown. The resting-state setting therefore leads to a challenging blind inverse problem, which is the focus of this work.
I-B Related Work
@@
The hemodynamic coupling is often modeled as a linear time-invariant system (LTI), in which the BOLD signal is expressed as a convolution of the underlying neural signals and the hemodynamic response function (HRF). In rsfMRI, many approaches avoid explicitly addressing the resulting blind inverse problem by assuming a known, spatially fixed HRF and then recover neural activity by solving the resulting (misspecified) linear inverse problem under a variety of spatio-temporal smoothness/sparsity priors [15, 16, 17]. In contrast, methods accommodating unknown hemodynamic couplings are less common: [18] use a template-shifted HRF and combine atlas-based low-rank and total-variation penalties for estimation; [19] use a two-stage procedure that first thresholds to detect activation and then estimates the HRF; [20] use a reparameterization approach based on the Wiener deconvolution and estimate the HRF using nonlinear least squares.
@@
A separate line of work uses nonlinear state-space models for coupling latent neural dynamics with more general, biophysically interpretable hemodynamic forward models [21], with inference procedures adapted to resting-state designs [22]. However, due to heavy computational costs, such approaches are generally limited to modeling a small number of brain regions [23], effectively imposing a strong limitation on spatial resolution. Moreover, for many practical acquisition protocols, several of the hemodynamic parameters in such models have questionable statistical identifiability [24].
@@
There is an extensive body of work addressing the closely related problem of blind image deblurring, where the goal is to recover both an unknown sharp image and the corresponding blur kernel from a noisy, blurred observation. Most approaches estimate image and kernel jointly, via alternating optimization [25] or deep learning based approaches [26]. In contrast, [27] outlines that, while the joint reconstruction problem is often severely ill-posed, the task can be decomposed into two more manageable subproblems: first marginalizing the image out of the data likelihood and estimating the blur by maximizing the resulting marginal likelihood, and then performing non-blind deblurring given the estimated kernel. While related, this class of blind image deblurring problems is structurally different from the blind hemodynamic inverse problem, as the former typically involves only a single unknown image-kernel pair, whereas the latter involves many coupled blind problems with complex underlying spatial dependencies in both signals and hemodynamics.
I-C Our Contribution
@@
In this work, we propose a novel methodology for estimating hemodynamic coupling in rsfMRI. In the spirit of [27], our estimator is based on a learned approximation to the marginal likelihood with respect to the latent neural signals, which enables us to decouple estimation of the forward hemodynamic coupling model from reconstruction of neural activity. The marginal likelihood is efficiently approximated using a conditional neural spline normalizing flow, trained using a biologically motivated neural–hemodynamic simulator model. This model is calibrated to real fMRI signals using a spectral moment–matching approach that guarantees the generative process produces physiologically plausible signals. Spatial regularization based on a latent Gaussian process prior is used to promote locally similar hemodynamic couplings across the cortical surface. We propose an estimation algorithm that leverages automatic differentiation of the learned likelihood and the sparsity structure of the Gaussian process prior to enable computational scalability to ultra high-resolution surface meshes. Uncertainty in the hemodynamic estimates is quantified via a double-bootstrap procedure, and automatic procedures for the selection of all model hyperparameters are proposed. Estimating the hemodynamic coupling enables the use of off-the-shelf methods to recover neural signals from BOLD data. Through extensive experiments using both synthetic and real data, we demonstrate substantial improvements over competing approaches in both hemodynamic estimation and downstream functional connectivity analysis.
II Models and Background
@@
This section outlines the models used for simulating neural activity and observed hemodynamic signals. Specifically, Section II-A introduces the neural activity model, Section II-B describes the signal-induced hemodynamic response model, and Section II-C proposes a statistical model for the observed rsfMRI data. Notably, the inversion method in Section III is largely simulator-agnostic, so alternative neural signal, hemodynamic, or noise models can be accommodated with minor modifications. Throughout, we will assume the fMRI time series are sampled at a constant repetition time $t_{r}>0$, with accompanying acquisition times $t_{m}=(m-1)t_{r}$, for m=1,,Mm=1,...,M, spanning the interval $0,(M-1)t_{r}]$.
II-A Modeling Neural Population Dynamics
@@
II-B Modeling the Hemodynamic Response
@@
II-C Statistical Model for Observed Data
For a set of cortical locations $\mathcal{X}:=\{x_{1},...,x_{V}\}\subset\Omega$, we gather noisy observations of the induced hemodynamic response at each of the acquisition times $t_{1},...,t_{M}$. The noise process contaminating the signals includes physiological and scanner contributions that may induce temporal autocorrelation [28]. For sufficiently small $t_{r}$, this autocorrelation can be modeled and removed during preprocessing, yielding an approximately white error process, which we assume here. The resulting Gaussian measurement noise model for the observed fMRI signals is then given by
@@ II-C Statistical Model for Observed Data
III Methodology
As outlined in Section I, direct joint estimation of the neural signals and hemodynamic parameters is often severely ill-posed and can lead to unstable inference. We instead marginalize over the neural signal generative process and base hemodynamic coupling estimation on the resulting (learned) marginal likelihood. To share information across space, we impose a Gaussian process prior on the hemodynamic parameter field and derive an estimator that is scalable to high-resolution cortical surface meshes. We then describe procedures for model hyperparameter selection and for quantifying estimation uncertainty. Finally, given the stage-one hemodynamic estimate, neural signal recovery reduces to a more stable non-blind inverse problem.
@@ III Methodology
III-A Learned Marginal Likelihood
@@
where the likelihood $p(y_{v}|s,\tilde{\theta},x_{v})$ is defined by (4), the point-wise signal prior $p(s|x_{v})$ is defined by (2), and $\mathcal{S}$ denotes the space of signals spanned by the generative model. From (4), it follows that $p(Y|\theta,\mathcal{X})=\prod_{v=1}^{V}p(y_{v}|\tilde{\theta}(x_{v}))$, hence it suffices to focus on the location-specific model in (5). Two major challenges arise. First, the marginalization integral in (5) does not admit a closed-form solution. Although a Monte Carlo approximation can be obtained by repeatedly simulating from (1) and (4), this is computationally prohibitive for large $V$, so a more scalable approximation is required. Second, $p(y_{v}|\tilde{\theta}(x_{v}))$ is typically high-dimensional, making it challenging to estimate. The remainder of this section outlines our approach to addressing these issues. For notational simplicity, we henceforth suppress the explicit dependence of (5) on $x$.
@@
In practice, we parameterize $T_{\psi}$ as a deep multilayer perceptron (MLP) with ReLU activations and $p_{\gamma}$ as a multilayer normalizing flow. More details on the network architecture and training can be found in Section IV-C.
III-B Prior Distribution
@@
On Euclidean domains, the Matérn processes are a popular choice for modeling spatial fields due to their flexibility and the closed-form of their covariance kernel. Following [30], such processes can be generalized to spatial signals located on the highly non-linear cortical surface $\Omega$ as the weak solutions to the following stochastic partial differential equation
@@
Assume the locations $x_{1},\ldots,x_{V}$ coincide with the vertices of the triangulated cortical surface $\Omega$, as is common in surface-based representation of fMRI. Using the linear finite element basis $\{\phi_{v}\}_{v=1}^{V}$ induced by this triangulation [31], define the inner product matrices with element-wise definitions
@@
III-C Point Estimation
We estimate the hemodynamic parameters of the model outlined in Section II by maximum a posteriori (MAP) estimation, combining the learned marginal likelihood from Section III-A with the spatial prior from Section III-B. Specifically, we aim to maximize the approximate posterior density
$p(\tilde{\boldsymbol{\theta}}|Y)$ $\approx p(\tilde{\boldsymbol{\theta}}|T(y_{1}),.,T(y_{V}))$ (10)
$\mathrel{\mkern-20.0mu\vbox{\halign{\hfil$#$\cr\propto\cr\kern 2.0pt\cr\sim\cr\kern-2.0pt\cr}}\mkern-20.0mu}\left[\prod_{v=1}^{V}p_{\widehat{\gamma}}(T_{\widehat{\psi}}(y_{v})|\tilde{\theta}_{v})\right]p(\tilde{\theta}_{1},.,\tilde{\theta}_{V}),$
@@ III-C Point Estimation
$\Bigl[-\nabla_{\text{vec}(\tilde{\boldsymbol{\theta}})}^{2}\!\log\prod_{v=1}^{V}p_{\widehat{\gamma}}\!\bigl(T_{\widehat{\psi}}(y_{v})\mid\tilde{\theta}_{v}\bigr)\Bigr]_{ii^{\prime}}=$ (16)
@@
III-D Hyperparameter Selection
III-D1 Calibration of Smoothness Prior
@@
$\log p(Y|\{\kappa_{j},\tau_{j}\}_{j})\approx\sum_{v=1}^{V}\log p_{\widehat{\gamma}}(T_{\widehat{\psi}}(y_{v})|\widehat{\tilde{\theta}}_{v})$ (17)
@@
III-D2 Calibration of Simulator
While $\theta_{j,\min},\theta_{j,\max}$ are assumed known a-priori, the models in Section II depend on several unknown simulator hyperparameters $\nu:=(\lambda_{\min},\lambda_{\max},a_{\min},a_{\max},\sigma^{2})$. To calibrate $\nu$ to our observed fMRI signals, we take a spectral moment-matching based approach. Specifically, let $p(\theta,s,\epsilon|\nu)$ denote the joint distribution induced by the simulator given $\nu$, and let $p^{*}(y)$ denote the true marginal distribution of the observed fMRI signals. We propose the following spectral moment-based loss for calibration

### v2

@@ Abstract
Functional magnetic resonance imaging (fMRI) provides an indirect measurement of neuronal activity via hemodynamic responses that vary across brain regions and individuals. Ignoring this hemodynamic variability can bias downstream connectivity estimates. Furthermore, the hemodynamic parameters themselves may serve as important imaging biomarkers. Estimating spatially varying hemodynamics from resting-state fMRI (rsfMRI) is therefore an important but challenging blind inverse problem, since both the latent neural activity and the hemodynamic coupling are unknown. In this work, we propose a methodology for inferring hemodynamic coupling on the cortical surface from rsfMRI. Our approach avoids the highly unstable joint recovery of neural activity and hemodynamics by marginalizing out the latent neural signal and basing inference on the resulting marginal likelihood. To enable scalable, high-resolution estimation, we employ a deep neural network combined with conditional normalizing flows to accurately approximate this intractable marginal likelihood, while enforcing spatial coherence through priors defined on the cortical surface that admit sparse representations. Uncertainty in the hemodynamic estimates is quantified via a double-bootstrap procedure. The proposed approach is extensively validated using synthetic data and real fMRI datasets, demonstrating clear improvements over current methods for hemodynamic estimation and downstream connectivity analysis.
keywords:
Hemodynamic coupling , hemodynamic inversion , deep learning , fMRI.
[…]
[label1]organization=Dept. of Statistics, University of South Carolina, addressline=1523 Greene St, city=Columbia, postcode=29225, state=SC, country=USA
[…]
[label2]organization=Dept. of Biostatistics, University of Washington, addressline=4333 Brooklyn Avenue NE, city=Seattle, postcode=98195, state=WA, country=USA
1 Introduction
1.1 Motivation
@@
Functional magnetic resonance imaging (fMRI) is a non-invasive imaging technique widely used to study neural activity. fMRI collects time-dependent blood oxygenation level–dependent (BOLD) signals, which reflect hemodynamic changes associated with neural activity (Ogawa et al., 1992). Accurate characterization of the hemodynamic coupling, i.e., the forward model that maps neural activity to BOLD signals, enables inference of underlying neural activity by solving the hemodynamic inverse problem (Buckner, 2003). However, this problem is complicated by the fact that the hemodynamic coupling is partially unknown and varies both across brain regions and between individuals (Loh et al., 2008; Bailes et al., 2023). Failure to account for the spatially varying hemodynamics can bias downstream estimates of functional connectivity (Lindquist et al., 2009; Rangaprakash et al., 2018, 2023). Additionally, the hemodynamic properties are themselves of independent interest as potential imaging biomarkers. Indeed, systematic hemodynamic alterations have been identified in a variety of neuropsychiatric conditions, including autism (Yan et al., 2018), schizophrenia and bipolar disorder (Yan et al., 2021), and obsessive compulsive disorder (Rangaprakash et al., 2021).
@@
There exists a substantial body of work on estimating hemodynamic coupling in task/event-based fMRI designs (Lange and Zeger, 2002; Wang et al., 2013; Zhang et al., 2012, 2013). In this setting, the neural signals are assumed to be known from the experimental design, making inversion considerably easier than in resting-state fMRI (rsfMRI), where both neural signals and hemodynamic response are unknown. The resting-state setting therefore leads to a challenging blind inverse problem, which is the focus of this work.
1.2 Related Work
@@
The hemodynamic coupling is often modeled as a linear time-invariant system (LTI), in which the BOLD signal is expressed as a convolution of the underlying neural signals and the hemodynamic response function (HRF). In rsfMRI, many approaches avoid explicitly addressing the resulting blind inverse problem by assuming a known, spatially fixed HRF and then recover neural activity by solving the resulting (misspecified) linear inverse problem under a variety of spatio-temporal smoothness/sparsity priors (Karahanoğlu et al., 2013; Caballero-Gaudes et al., 2019; Uruñuela et al., 2024). In contrast, methods accommodating unknown hemodynamic couplings are less common: Cherkaoui et al. (2021) use a template-shifted HRF and combine atlas-based low-rank and total-variation penalties for estimation; Wu et al. (2013) use a two-stage procedure that first thresholds to detect activation and then estimates the HRF; Singh et al. (2020) use a reparameterization approach based on the Wiener deconvolution and estimate the HRF using nonlinear least squares.
@@
A separate line of work uses nonlinear state-space models for coupling latent neural dynamics with more general, biophysically interpretable hemodynamic forward models (Buxton et al., 1998), with inference procedures adapted to resting-state designs (Havlicek et al., 2011). However, due to heavy computational costs, such approaches are generally limited to modeling a small number of brain regions (Cao et al., 2019), effectively imposing a strong limitation on spatial resolution. Moreover, for many practical acquisition protocols, several of the hemodynamic parameters in such models have questionable statistical identifiability (Zayane and Laleg-Kirati, 2015).
@@
There is an extensive body of work addressing the closely related problem of blind image deblurring, where the goal is to recover both an unknown sharp image and the corresponding blur kernel from a noisy, blurred observation. Most approaches estimate image and kernel jointly, via alternating optimization (Chan and Wong, 1998) or deep learning based approaches (Gossard and Weiss, 2024). In contrast, Wipf and Zhang (2014) outlines that, while the joint reconstruction problem is often severely ill-posed, the task can be decomposed into two more manageable subproblems: first marginalizing the image out of the data likelihood and estimating the blur by maximizing the resulting marginal likelihood, and then performing non-blind deblurring given the estimated kernel. While related, this class of blind image deblurring problems is structurally different from the blind hemodynamic inverse problem, as the former typically involves only a single unknown image-kernel pair, whereas the latter involves many coupled blind problems with complex underlying spatial dependencies in both signals and hemodynamics.
1.3 Our Contribution
@@
In this work, we propose a novel methodology for estimating hemodynamic coupling in rsfMRI. In the spirit of Wipf and Zhang (2014), our estimator is based on a learned approximation to the marginal likelihood with respect to the latent neural signals, which enables us to decouple estimation of the forward hemodynamic coupling model from reconstruction of neural activity. The marginal likelihood is efficiently approximated using a conditional neural spline normalizing flow, trained using a biologically motivated neural–hemodynamic simulator model. This model is calibrated to real fMRI signals using a spectral moment–matching approach that guarantees the generative process produces physiologically plausible signals. Spatial regularization based on a latent Gaussian process prior is used to promote locally similar hemodynamic couplings across the cortical surface. We propose an estimation algorithm that leverages automatic differentiation of the learned likelihood and the sparsity structure of the Gaussian process prior to enable computational scalability to ultra high-resolution surface meshes. Uncertainty in the hemodynamic estimates is quantified via a double-bootstrap procedure, and automatic procedures for the selection of all model hyperparameters are proposed. Estimating the hemodynamic coupling enables the use of off-the-shelf methods to recover neural signals from BOLD data. Through extensive experiments using both synthetic and real data, we demonstrate substantial improvements over competing approaches in both hemodynamic estimation and downstream functional connectivity analysis.
2 Models and Background
@@
This section outlines the models used for simulating neural activity and observed hemodynamic signals. Specifically, Section 2.1 introduces the neural activity model, Section 2.2 describes the signal-induced hemodynamic response model, and Section 2.3 proposes a statistical model for the observed rsfMRI data. Figure 1 illustrates the generative process. Notably, the inversion method in Section 3 is largely simulator-agnostic, so alternative neural signal, hemodynamic, or noise models can be accommodated with minor modifications. Throughout, we will assume the fMRI time series are sampled at a constant repetition time $t_{r}>0$, with accompanying acquisition times $t_{m}=(m-1)t_{r}$, for m=1,,Mm=1,...,M, spanning the interval $0,(M-1)t_{r}]$.
Figure 1: Simulation from the data generative model. (Left) Neural activity sampled from (2). (Middle) Hemodynamic response obtained by convolving the neural activity with the HRF (20) under forward model (3). (Right) Simulated BOLD measurements generated according to (4) ($t_{r}=0.72\,\mathrm{s}$).
2.1 Modeling Neural Population Dynamics
@@
2.2 Modeling the Hemodynamic Response
@@
2.3 Statistical Model for Observed Data
For a set of cortical locations $\mathcal{X}:=\{x_{1},...,x_{V}\}\subset\Omega$, we gather noisy observations of the induced hemodynamic response at each of the acquisition times $t_{1},...,t_{M}$. The noise process contaminating the signals includes physiological and scanner contributions that may induce temporal autocorrelation (Lindquist, 2008). For sufficiently small $t_{r}$, this autocorrelation can be modeled and removed during preprocessing, yielding an approximately white error process, which we assume here. The resulting Gaussian measurement noise model for the observed fMRI signals is then given by
@@ 2.3 Statistical Model for Observed Data
3 Methodology
As outlined in Section 1, direct joint estimation of the neural signals and hemodynamic parameters is often severely ill-posed and can lead to unstable inference. We instead marginalize over the neural signal generative process and base hemodynamic coupling estimation on the resulting (learned) marginal likelihood. To share information across space, we impose a Gaussian process prior on the hemodynamic parameter field and derive an estimator that is scalable to high-resolution cortical surface meshes. We then describe procedures for model hyperparameter selection and for quantifying estimation uncertainty. Finally, given the stage-one hemodynamic estimate, neural signal recovery reduces to a more stable non-blind inverse problem.
@@ 3 Methodology
3.1 Learned Marginal Likelihood
@@
where the likelihood $p(y_{v}|s,\tilde{\theta},x_{v})$ is defined by (4), the point-wise signal prior $p(s|x_{v})$ is defined by (2), and $\mathcal{S}$ denotes the space of signals spanned by the generative model. For tractability, we ignore the complex spatial dependence among neural signals and instead we approximate the joint prior over $\mathcal{X}$ by a product of the marginal priors outlined in Section 2.1. It follows that $p(Y|\theta,\mathcal{X})\approx\prod_{v=1}^{V}p(y_{v}|\tilde{\theta}(x_{v}))$, hence it suffices to focus on the location-specific model in (5). Two major challenges arise. First, the marginalization integral in (5) does not admit a closed-form solution. Although a Monte Carlo approximation can be obtained by repeatedly simulating from (1) and (4), this is computationally prohibitive for large $V$, so a more scalable approximation is required. Second, $p(y_{v}|\tilde{\theta}(x_{v}))$ is typically high-dimensional, making it challenging to estimate. The remainder of this section outlines our approach to addressing these issues. For notational simplicity, we henceforth suppress the explicit dependence of (5) on $x$.
@@
In practice, we parameterize $T_{\psi}$ as a deep multilayer perceptron (MLP) with ReLU activations and $p_{\gamma}$ as a multilayer normalizing flow. More details on the network architecture and training can be found in Section 4.3.
3.2 Prior Distribution
@@
On Euclidean domains, the Matérn processes are a popular choice for modeling spatial fields due to their flexibility and the closed-form of their covariance kernel. Following Lindgren et al. (2011), such processes can be generalized to spatial signals located on the highly non-linear cortical surface $\Omega$ as the weak solutions to the following stochastic partial differential equation
@@
Assume the locations $x_{1},\ldots,x_{V}$ coincide with the vertices of the triangulated cortical surface $\Omega$, as is common in surface-based representation of fMRI. Using the linear finite element basis $\{\phi_{v}\}_{v=1}^{V}$ induced by this triangulation (Lai and Schumaker, 2007), define the inner product matrices with element-wise definitions
@@
3.3 Point Estimation
We estimate the hemodynamic parameters of the model outlined in Section 2 by maximum a posteriori (MAP) estimation, combining the learned marginal likelihood from Section 3.1 with the spatial prior from Section 3.2. Specifically, we aim to maximize the approximate posterior density
$p(\tilde{\boldsymbol{\theta}}|Y)$ $\approx p(\tilde{\boldsymbol{\theta}}|T(y_{1}),.,T(y_{V}))\mathrel{\mkern-20.0mu\vbox{\halign{\hfil$#$\cr\propto\cr\kern 2.0pt\cr\sim\cr\kern-2.0pt\cr}}\mkern-20.0mu}\left[\prod_{v=1}^{V}p_{\widehat{\gamma}}(T_{\widehat{\psi}}(y_{v})|\tilde{\theta}_{v})\right]p(\tilde{\theta}_{1},.,\tilde{\theta}_{V}),$ (10)
@@ 3.3 Point Estimation
3.4 Hyperparameter Selection
3.4.1 Calibration of Smoothness Prior
@@
[…]
@@
3.4.2 Calibration of Simulator
While $\theta_{j,\min},\theta_{j,\max}$ are assumed known a-priori, the models in Section 2 depend on several unknown simulator hyperparameters $\nu:=(\lambda_{\min},\lambda_{\max},a_{\min},a_{\max},\sigma^{2})$. To calibrate $\nu$ to our observed fMRI signals, we take a spectral moment-matching based approach. Specifically, let $p(\theta,s,\epsilon|\nu)$ denote the joint distribution induced by the simulator given $\nu$, and let $p^{*}(y)$ denote the true marginal distribution of the observed fMRI signals. We propose the following spectral moment-based loss for calibration